lass: Sta For each of the following situations, find the critical value(s) for z or t. a) Ho: ρ 0.2 vs. HA: ρ #0.2 at α-0.05 b) Ho: ρ 0.5 vs. HA: ρ > 0.5 at α-0.05 c) Ho: μ-40 vs. HA: μ#40 at α-0.05; n.50 d)Ho: ρ#0.2 vs. HA, p>O.2 at α 0.10; n-340 e) Ho: μ-80 vs. HA: μ < 80 at α :0.05; n-1000 ork
Answer #1

a)

Test is two tailed so critical values of z are: -1.96 and 1.96

Excel function used: "=NORMSINV(0.025)" and "=NORMSINV(0.975)"

b)

Test is right tailed so critical value of z is: 1.645

Excel function used: "=NORMSINV(0.95)"

c)

Degree of freedom:

df =n-1=49

The critical value of t are: -2.01 and 2.01

Excel function used: "=TINV(0.05,49)"

d)

Since test is for proportion so z critical value will be used.

Test is right tailed so critical value of z is: 1.28

Excel function used: "=NORMSINV(0.90)"

e)

Here sample size very large sowe can use z-critical value instead of t-critical values.

Test is left tailed so critical value of z is: -1.645

Excel function used: "=NORMSINV(0.05)"